OpenStudy (anonymous):

Solve the following for θ, where θ is greater than or equal to 0 and less than 2π: 2cos^2θ-5cosθ=3 I factored it out to (cos^2θ-3)(2cos^2θ+1)=0 but got answers of cosθ=(√3) and cosθ=(√-1/2) which doesn't make sense so I think i'm doing it all wrong.

OpenStudy (campbell_st):

well your factorisation is slightly wrong try \[(\cos(\theta) -3)(2\cos(\theta) + 1) = 0\] so the solutions are \[\cos(\theta) = 3, \frac{-1}{2}\] the\[\cos(\theta) = 3\] solution doesn't exist so ignore it as a solution just find the values where \[\cos(\theta) = \frac{-1}{2}\] in the given domian

OpenStudy (anonymous):

Ahh thank you, that makes much more sense! I got 2π/3 and 4π/3 :)

OpenStudy (campbell_st):

yep... thats it..